# Optimizing Distributed Training on Frontier for Large Language Models

### 20 Dec 2023 v1

@@ Abstract
Large language models (LLM) are showing tremendous success as foundation models, and many downstream applications benefit from fine-tuning. Prior works on loss scaling have demonstrated that the larger LLMs perform better than their smaller counterparts. However, training LLMs with billions of parameters requires considerable computational resources; to train a one trillion GPT-style model on 20 trillion tokens, we need to perform 120 million exaflops. Frontier is the world’s first and fastest exascale supercomputer for open science and is equipped with 75264 MI250X GPUs. This work explores efficient distributed strategies such as tensor parallelism, pipeline parallelism, and sharded data parallelism to train a trillion-parameter model on the Frontier exascale supercomputer. We analyze these distributed training techniques and associated parameters individually to decide which techniques to use and what associated parameters to select for a particular technique. We perform hyperparameter tuning on these techniques to understand their complex interplay. Combined with these two tuning efforts, we have found optimal strategies to train three models of size 22B, 175B, and 1T parameters with $38.38\%$ , $36.14\%$ , and $31.96\%$ achieved throughput. For training the 175B parameter model and 1T model, we have achieved $100\%$ weak scaling efficiency and $89\%$ and $87\%$ strong scaling efficiency, respectively. Our work presents a set of strategies for distributed training of LLMs through experimental findings and hyperparameter tuning.
@@ III-B1 Impact of Number of Pipeline Stages
IV Distribution Parameter Tuning Using DeepHyper
@@ VI Conclusion and Discussion
Training an LLM with Billions to Trillions of parameters is a challenging task since we have to orchestrate overcoming the GPU memory wall, minimizing communication latency, and building a software stack with state-of-the-art distribution algorithms. A Trillion parameter model requires a minimum of 14 Terabytes of memory, while an MI250X GPU has only 64 Gigabytes. So, to overcome the memory wall problem, we have explored a combination of model parallelism strategies. However, model parallel training requires communication within groups of GPUs either sharing the parts of the same tensor (Tensor parallelism) or the groups of GPUs hosting neighboring components (pipeline parallelism). We also need to utilize data parallelism to consume a large number of tokens simultaneously and use a larger number of computing resources to achieve faster time to solution. All of these model parallelism and data parallelism incur communication at various frequencies and of various volumes. The communication latency can significantly increase the training time and reduce the training performance regarding GPU throughput.

### 21 Dec 2023 v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable success as foundational models, benefiting various downstream applications through fine-tuning. Recent studies on loss scaling have demonstrated the superior performance of larger LLMs compared to their smaller counterparts. Nevertheless, training LLMs with billions of parameters poses significant challenges and requires considerable computational resources. For example, training a one trillion parameter GPT-style model on 20 trillion tokens requires a staggering 120 million exaflops of computation. This research explores efficient distributed training strategies to extract this computation from Frontier, the world’s first exascale supercomputer dedicated to open science. We enable and investigate various model and data parallel training techniques, such as tensor parallelism, pipeline parallelism, and sharded data parallelism, to facilitate training a trillion-parameter model on Frontier. We empirically assess these techniques and their associated parameters to determine their impact on memory footprint, communication latency, and GPU’s computational efficiency. We analyze the complex interplay among these techniques and find a strategy to combine them to achieve high throughput through hyperparameter tuning. We have identified efficient strategies for training large LLMs of varying sizes through empirical analysis and hyperparameter tuning. For 22 Billion, 175 Billion, and 1 Trillion parameters, we achieved GPU throughputs of $38.38\%$, $36.14\%$, and $31.96\%$, respectively. For the training of the 175 Billion parameter model and the 1 Trillion parameter model, we achieved $100\%$ weak scaling efficiency on 1024 and 3072 MI250X GPUs, respectively. We also achieved strong scaling efficiencies of $89\%$ and $87\%$ for these two models.
@@ III-B1 Impact of Number of Pipeline Stages
IV Hyperparameter Tuning Using DeepHyper
@@ VI Conclusion and Discussion
Training an LLM with Billions to Trillions of parameters is a challenging task since we have to orchestrate overcoming the GPU memory wall, minimizing communication latency, and building a software stack with state-of-the-art distribution algorithms. A Trillion parameter model requires a minimum of 14 Terabytes of memory, while an MI250X GPU has only 64 Gigabytes. So, to overcome the memory wall problem, we have explored a combination of model parallelism strategies. However, model parallel training requires communication within groups of GPUs either sharing the parts of the same tensor (Tensor parallelism) or the groups of GPUs hosting neighboring components (pipeline parallelism). We also needed to utilize data parallelism to consume a large number of tokens simultaneously and use a larger number of computing resources to achieve faster time to solution. All of these model parallelism and data parallelism incur communication at various frequencies and of various volumes. The communication latency can significantly increase the training time and reduce the training performance regarding GPU throughput.
@@
Acknowledgements
This research was sponsored by and used resources of the Oak Ridge Leadership Computing Facility (OLCF), which is a DOE Office of Science User Facility at the Oak Ridge National Laboratory supported by the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.